OpenStudy (anonymous):

11/2+3x=-5 by the power of 2 /2 I keep getting the same exact answer of 7/3 but my book says -6 . I have checked this problem 10 times and it keeps coming out to the same outcome could someone elaborate on what i am doing wrong. The book says the answer is -6 how is that possible?

OpenStudy (anonymous):

Is the equation \[\frac{ 11 }{ 2+3x }= -5\]?

OpenStudy (anonymous):

No its 11/2 +3x= -5 to the power of 2 / by 2

OpenStudy (anonymous):

\[\frac{ 11 }{ 2 }+3x=\frac{ -5^2 }{ 2}\]

OpenStudy (anonymous):

11/2+3x=-25/2

OpenStudy (anonymous):

i keep getting 7/3

OpenStudy (anonymous):

3x=-36/2

OpenStudy (anonymous):

x=-6

OpenStudy (anonymous):

oh wtf

OpenStudy (anonymous):

do that agin

OpenStudy (anonymous):

-5^2/2=-25/2

OpenStudy (anonymous):

but (-5)^2/2=25/2, which is the part where you had a problem with

OpenStudy (anonymous):

thats impossible you have 11/2+3x= 25/2

OpenStudy (anonymous):

how did you get to 36

OpenStudy (anonymous):

11/2+3x= -25/2

OpenStudy (anonymous):

yes if you divided you get a mixed number

OpenStudy (anonymous):

but that has no effect on the product so how did you do it

OpenStudy (anonymous):

3x= -25/2-11/2

OpenStudy (anonymous):

3x=-36/2

OpenStudy (anonymous):

3x=-18

OpenStudy (anonymous):

woah how did you move it

OpenStudy (anonymous):

11/2+3x-11/2= -25/2-11/2

OpenStudy (anonymous):

5 times 5 is a positive not a negative

OpenStudy (anonymous):

-5^2=-25

OpenStudy (anonymous):

without a bracket, its still negative

OpenStudy (anonymous):

wtf

OpenStudy (anonymous):

(-5)^2=25

OpenStudy (anonymous):

i never knew this

OpenStudy (anonymous):

now it would make sense

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

-5 squared without a bracket is 25 isolated by itself

OpenStudy (anonymous):

thats impossible

OpenStudy (anonymous):

its a whole fraction you're dealing with \[\frac{ -5^2 }{ 2}=\frac{ 5^2 }{ -2}\]

OpenStudy (anonymous):

oh pellet

OpenStudy (anonymous):

oh pellet you mean the whole fraction was negative so the numerator can't be negative by itself?

OpenStudy (anonymous):

thats very misleading then sense the book put -5 negative in itself

OpenStudy (anonymous):

yeah, think of it like this \[-(\frac{ 5^2 }{ 2})\]

OpenStudy (anonymous):

Oh ok let me preform the problem

OpenStudy (anonymous):

Beautiful